Light from a source of wavelength 475 nm passes through a single slit and falls on a screen 2.5 m away. If the slit width is 2.00 μm, how many maxima occur, including the central maximum?
1
Expert's answer
2019-04-10T09:44:41-0400
"d \sin \theta = n\lambda"

"n_{max}=\frac{d}{\lambda}=\frac{2}{0.475}=4"

Thus, there are 4 minima from each side of  the central maximum. So, there are


"2(4-1)+1=7"

maxima including the central maximum.